# Balancing Speciality and Versatility: a Coarse to Fine Framework for Supervised Fine-tuning Large Language Model

### v2

@@ Abstract
Aligned Large Language Models (LLMs) showcase remarkable versatility, capable of handling diverse real-world tasks. Meanwhile, aligned LLMs are also expected to exhibit speciality, excelling in specific applications. However, fine-tuning with extra data, a common practice to gain speciality, often leads to catastrophic forgetting (CF) of previously acquired versatility, hindering the model’s performance across diverse tasks. In response to this challenge, we propose CoFiTune, a coarse to fine framework in an attempt to strike the balance between speciality and versatility. At the coarse-grained level, an empirical tree-search algorithm is utilized to pinpoint and update specific modules that are crucial for speciality, while keeping other parameters frozen; at the fine-grained level, a soft-masking mechanism regulates the update to the LLMs, mitigating the CF issue without harming speciality. In an overall evaluation of both speciality and versatility, CoFiTune consistently outperforms baseline methods across diverse tasks and model scales. When compared to the full-parameter SFT, CoFiTune offers an average versatility improvement of 14%, while only incurring a marginal loss in speciality. Lastly, based on further analysis, we provide a speculative insight into the information forwarding process in LLMs, which helps explain the effectiveness of the proposed method. The code is available at https://github.com/rattlesnakey/CoFiTune.
@@
Hengyuan Zhang1†thanks: †Work was done during internship at SenseTime Research, Yanru Wu1, Dawei Li3, Zacc Yang2, Rui Zhao2, Yong Jiang1*, Fei Tan2*\$thanks: * Corresponding authorthanks: \$ Project leader 1Tsinghua University  2SenseTime Research  3University of California, San Diego {zhang-hy22,wu-yr21}@mails.tsinghua.edu.cn  dal034@ucsd.edu {yangzacc,zhaorui,tanfei}@sensetime.com
@@ 1 Introduction
However, during the fine-tuning process, the modification of model parameters often leads to catastrophic forgetting (CF), causing a noticeable loss of versatility (Lin et al., 2023b). This loss adversely affects the performance of fine-tuned models across various real-world tasks (Cheng et al., 2023; Dong et al., 2023), propelling several works to investigate and contribute solutions to the CF in LLM versatility (Lin et al., 2023b; Wang et al., 2023c).
As a relatively new problem, the CF in LLMs remains under-explored. We categorize the existing studies into regularization-based, weight-based, and architecture-based methods [^1] See Appendix A.5 for more details about these methods.. Regularization-based methods (Lin et al., 2023b; Smith et al., 2023) add extra terms into loss function to penalize parameter changes. Weight-based methods (Wortsman et al., 2022; Ke et al., 2022) introduce weight coefficients for parameters to regulate their updates. Architecture-based methods (Wang et al., 2023d; Razdaibiedina et al., 2023) design and exclusively fine-tune extra modules outside the original model. However, these methods still have limitations: 1) Regularization- and weight-based methods involve fine-tuning all parameters, which poses a significant challenge to preventing CF in versatility, especially as the training iteration increases, CF becomes severer (Luo et al., 2023b). 2) Architecture-based methods only update the inserted parameters, which blocks the learning of speciality, particularly on challenging tasks (Lin et al., 2023b).
These limitations raise a key question: How can the model gain speciality while preserving the versatility, thereby enhancing its overall abilities? (shown in Fig. 1). Recent research highlights two key findings: the presence of redundant parameters in the model (Bansal et al., 2023; Sun et al., 2023) and the distinct role played by different modules and layers (Geva et al., 2021; Meng et al., 2022a). These suggest that it is highly feasible to enhance speciality by updating specific parameters within a defined layer range while keeping the remainder frozen to maintain versatility. [^2] The results in Sec. 4.3 and 4.4, along with the insights 1, 3, and 4, further support our speculation.
@@ 2 Related Work
Previous works demonstrate the presence of substantial redundant parameters in Transformer-based models (An et al., 2020; Xia et al., 2022; Bansal et al., 2023). Therefore, it’s crucial to identify the key components in it and accurately comprehend their internal mechanisms. Recent endeavors try to analyze the layers and modules in it. Elhage et al. (2021) developed mathematical tools to unveil an information-copying behavior within the attention module. Geva et al. (2021) analyzed the feed-forward network (FFN), considering the up and down projection matrices as the key and value of the memories. Mirzadeh et al. (2024) demonstrated that SwiGLU, a common trainable activation function in the FFN of LLMs, can be replaced with non-trainable ReLU to reduce computation without affecting performance. Based on the aforementioned insights, in this work, we explore the key components, i.e., attention and FFN [^4] For simplicity, starting from this section, FFN only refers to the combination of up & down projection. modules in certain layer range that are crucial for gaining speciality and preserving versatility.
@@ 4 Experiment
In this section, we conduct extensive experiments to answer the following questions: 1) RQ1: Training which layers and modules can optimize the trade-off between speciality and versatility? 2) RQ2: How does our CoFiTune framework compared to the baseline methods? 3) RQ3: Can the Fine-SoftMask mechanism further alleviate the CF issue in versatility? 4) RQ4: Can our CoFiTune be seamlessly applied to models across different scales and families?
@@ 4.1 Datasets and Experiment Settings
Our setting emphasizes both speciality and versatility. Specifically, for LLM’s speciality, we select tasks with considerable complexity in two primary categories: 1) Knowledge Intensive Tasks: These tasks require the model to generate responses to questions tailored to specific domain. We create the Finance and Law instruction format datasets based on FiQA (Wang et al., 2023b) and LlamaLawyer (Huang et al., 2023); 2) Ability Intensive Tasks: These tasks aim to test LLM’s ability to handle challenging and crucial tasks in real-world scenarios. We create the mathematical problem solving (Math) and Context-Aware Generation (CAG [^8] CAG task requires the model to generate answers flexibly based on the given context (See Sec. A.6.2 for more details).) based on MGSM8k (Chen et al., 2023a) and DuReader (He et al., 2018).
@@ 4.4 Further Analysis
Furthermore, combining the hints from Geva et al. (2021), Meng et al. (2022b), and Chuang et al. (2023) with our findings in this work, we speculate the forward process in Fig. 5: 1) MHA modules in Versatility Area aggregate information, enriching input for corresponding FFN. FFN modules utilize this enriched input information to extract relevant knowledge, passing it through residual connections to Optimal Speciality Gaining Area; 2) Optimal Speciality Gaining Area handles both versatility and speciality information, forwarding it to Augmenting Area. 3) Processed information from the previous two areas is forwarded to Augmenting Area through residual connections, focusing on information enhancement for final prediction.
@@ A.6 Statistics and Descriptions of Datasets
Currently, LLM still falls short in certain areas (Li et al., 2023a; Lu et al., 2023a; Zhang et al., 2024a, b; Li et al., 2024), we mainly select two types of tasks to evaluate its speciality. The overall statistics of the selected task datasets are shown in Table 7, and the detailed descriptions are as follows:
@@ A.10.2 Speciality Score
For Finance, Law, and CAG tasks, we follow Zhang et al. (2022a); Kong et al. (2022b, a); Zhang et al. (2023) to utilize the automatic metrics to evaluate the quality of generated text , i.e., BERTScore [^23] https://github.com/Tiiiger/bert_score, BLEU [^24] https://www.nltk.org/api/nltk.translate.bleu_score.html, and Rouge [^25] https://github.com/pltrdy/rouge to evaluate their speciality performance. The Rouge metric comprises Rouge-1, Rouge-2, and Rouge-L. We initially average the results of Rouge-1, Rouge-2, and Rouge-L to obtain the overall Rouge result. Subsequently, we further average the results of BERTScore, BLEU, and Rouge to derive the final speciality score. It’s worth noting that we employ the F1 score in both Rouge and BERTScore.

### v5

@@ Abstract
Aligned Large Language Models (LLMs) showcase remarkable versatility, capable of handling diverse real-world tasks. Meanwhile, aligned LLMs are also expected to exhibit speciality, excelling in specific applications. However, fine-tuning with extra data, a common practice to gain speciality, often leads to catastrophic forgetting (CF) of previously acquired versatility, hindering the model’s performance across diverse tasks. In response to this challenge, we propose CoFiTune, a coarse to fine framework in an attempt to strike the balance between speciality and versatility. At the coarse-grained level, an empirical tree-search algorithm is utilized to pinpoint and update specific modules that are crucial for speciality, while keeping other parameters frozen; at the fine-grained level, a soft-masking mechanism regulates the update to the LLMs, mitigating the CF issue without compromising speciality. In an overall evaluation of both speciality and versatility, CoFiTune consistently outperforms baseline methods across diverse tasks and model scales. When compared to the full-parameter SFT, CoFiTune offers an average versatility improvement of 14%, while only incurring a marginal loss in speciality. Lastly, based on further analysis, we provide a speculative insight into the information forwarding process in LLMs, which helps explain the effectiveness of the proposed method. The code is available at https://github.com/rattlesnakey/CoFiTune.
@@
Hengyuan Zhang1†thanks: †Work was done during internship at SenseTime Research, Yanru Wu1†, Dawei Li3, Sak Yang4, Rui Zhao2, Yong Jiang1*, Fei Tan2*\$thanks: * Corresponding authorthanks: \$ Project lead 1Tsinghua University  2SenseTime Research 3University of California, San Diego  4Independent Researcher {zhang-hy22,wu-yr21}@mails.tsinghua.edu.cn  dal034@ucsd.edu sakyang@outlook.com  {zhaorui,tanfei}@sensetime.com
@@ 1 Introduction
However, during the fine-tuning process, the modification of model parameters often leads to catastrophic forgetting (CF), thereby causing a noticeable loss of versatility (Lin et al., 2023b). This loss adversely affects the performance of fine-tuned models across various real-world tasks (Cheng et al., 2023; Dong et al., 2023), propelling several works to investigate and contribute solutions to the CF in LLM versatility (Lin et al., 2023b; Wang et al., 2023c).
As a relatively new problem, the CF in LLMs remains under-explored. We categorize the existing studies into regularization-based, weight-based, and architecture-based methods [^1] See Appendix A.5 for more details about these methods.. Regularization-based methods (Lin et al., 2023b; Smith et al., 2023) add extra terms into loss function to penalize parameter changes. Weight-based methods (Wortsman et al., 2022; Ke et al., 2022) introduce weight coefficients for parameters to regulate their updates. Architecture-based methods (Wang et al., 2023d; Razdaibiedina et al., 2023) design and exclusively fine-tune extra modules outside the original model. However, these methods still have limitations: 1) Regularization- and weight-based methods involve fine-tuning all parameters, which poses a significant challenge to preventing CF in versatility. Particullary, CF deteriorates as the training iteration proceeds (Luo et al., 2023b). 2) Architecture-based methods only update the inserted parameters, which blocks the learning of speciality, particularly on challenging tasks (Lin et al., 2023b).
These limitations raise a key question: How can the model gain speciality while preserving the versatility, thereby enhancing its overall abilities? (shown in Fig. 1). Recent research threads highlight two pivotal findings: the presence of redundant parameters in the model (Bansal et al., 2023; Sun et al., 2023) and the distinct role played by different modules and layers (Geva et al., 2021; Meng et al., 2022a). These suggest that it is highly feasible to enhance speciality by updating specific parameters within a defined layer range while keeping the remainder frozen to maintain versatility. [^2] The results in Sec. 4.3 and 4.4, along with the insights 1, 3, and 4, further support our speculation.
@@ 2 Related Work
Previous works demonstrate the presence of substantial redundant parameters in Transformer-based models (An et al., 2020; Xia et al., 2022; Bansal et al., 2023). Therefore, it’s crucial to pinpoint the key components and precisely comprehend their underlying mechanisms. Recent endeavors try to analyze the involved layers and modules. Elhage et al. (2021) developed mathematical tools to unveil an information-copying behavior within the attention module. Geva et al. (2021) analyzed the feed-forward network (FFN), considering the up and down projection matrices as the key and value of the memories. Mirzadeh et al. (2024) demonstrated that SwiGLU, a common trainable activation function in the FFN of LLMs, can be replaced with non-trainable ReLU to reduce computation without affecting performance. In light of the aforementioned insights, in this work, we explore the key components, i.e., attention and FFN [^4] For simplicity, starting from this section, FFN only refers to the combination of up & down projection. modules in certain layer range that are integral to gaining speciality and preserving versatility.
@@ 4 Experiment
In this section, we conduct extensive experiments to answer the following questions: 1) RQ1: Training which layers and modules can optimize the trade-off between speciality and versatility? 2) RQ2: How does our CoFiTune framework compared to the baseline methods? 3) RQ3: Can the Fine-SoftMask mechanism further alleviate the CF issue in versatility? 4) RQ4: Can our CoFiTune be seamlessly applied to models across different model scales and families?
@@ 4.1 Datasets and Experiment Settings
Our setting emphasizes both speciality and versatility. Specifically, for LLM’s speciality, we select tasks with considerable complexity in two primary categories: 1) Knowledge Intensive Tasks: These tasks require the model to generate responses to questions tailored for specific domain. We create the Finance and Law instruction format datasets based on FiQA (Wang et al., 2023b) and LlamaLawyer (Huang et al., 2023); 2) Ability Intensive Tasks: These tasks aim to test LLM’s ability to handle challenging and crucial tasks in real-world scenarios. We create the mathematical problem solving (Math) and Context-Aware Generation (CAG [^8] CAG task requires the model to generate answers flexibly based on the given context (See Sec. A.6.2 for more details).) based on MGSM8k (Chen et al., 2023a) and DuReader (He et al., 2018).
@@ 4.4 Further Analysis
Furthermore, combining the hints from Geva et al. (2021), Meng et al. (2022b), Allen-Zhu and Li (2023b), Allen-Zhu and Li (2023a), and Chuang et al. (2023) with our findings in this work, we speculate the forward process in Fig. 5: 1) MHA modules in Versatility Area aggregate information, enriching input for corresponding FFN. FFN modules utilize this enriched input information to extract relevant knowledge, passing it through residual connections to Optimal Speciality Gaining Area; 2) Optimal Speciality Gaining Area handles both versatility and speciality information, forwarding it to Augmenting Area. 3) Processed information from the previous two areas is forwarded to Augmenting Area through residual connections, focusing on information enhancement for final prediction.
@@ A.6 Statistics and Descriptions of Datasets
Currently, LLM still falls short in certain areas (Li et al., 2023a; Lu et al., 2023a), we mainly select two types of tasks to evaluate its speciality. The overall statistics of the selected task datasets are shown in Table 7, and the detailed descriptions are as follows:
@@ A.10.2 Speciality Score
For Finance, Law, and CAG tasks, we follow Tan et al. (2020, 2021); Zhang et al. (2022a); Kong et al. (2022b, a); Zhang et al. (2023) to utilize the automatic metrics to evaluate the quality of generated text , i.e., BERTScore [^23] https://github.com/Tiiiger/bert_score, BLEU [^24] https://www.nltk.org/api/nltk.translate.bleu_score.html, and Rouge [^25] https://github.com/pltrdy/rouge to evaluate their speciality performance. The Rouge metric comprises Rouge-1, Rouge-2, and Rouge-L. We initially average the results of Rouge-1, Rouge-2, and Rouge-L to obtain the overall Rouge result. Subsequently, we further average the results of BERTScore, BLEU, and Rouge to derive the final speciality score. It’s worth noting that we employ the F1 score in both Rouge and BERTScore.